OpenStudy (anonymous):

What is the mathematical equation of the sum of two consecutive integers is equals to 25?

OpenStudy (p0sitr0n):

first:n second : n+1 first+second=N+n+1=2n+1=25 2n=24 n=12 first:12 second:13

OpenStudy (ahsome):

Consecutive numbers are numbers that differ between itself by 1 Let the first number be \(x\) Since they are consecutive, the second number is \(x+1\) \[x+x+1=25\]

OpenStudy (ahsome):

If you go through the equation, you can find that \(x=12\). Then, you can add one and find that the second number equals \(13\)

OpenStudy (anonymous):

what if it is product not sum?